OpenStudy (anonymous):

write this ratio as a fraction in lowest terms? 100 to 90

OpenStudy (anonymous):

to write this as a fraction just write is as a/b

OpenStudy (anonymous):

so in this case, it's \[\frac{ 100 }{ 90 }\]

OpenStudy (anonymous):

now when simplifying fractions, you have to do the same operation, usually division, to the top and bottom of the fraction. So think of a number that will divide into 100 and 90 without a remainder.\[\frac{ 100 }{ 90 }\frac{ \div10 }{ \div10}= \frac{ 10 }{ 9 }\]

OpenStudy (anonymous):

there's your answer

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem